OpenStudy (anonymous):

MEDALS AND FAN demonstrate the relationship between simple interest and linear growth

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

The formula for simple interest is B = A (1 + r*X), where B = Balance, A = original balance, X = number of periods of interest, and r = interest rate. B = A (1 + r*X) is a linear function. Thus, simple interest grows linearly.

OpenStudy (anonymous):

define compound interest as a real-life example of exponential growth

OpenStudy (anonymous):

@mathmate